A 60.0 mL sample of 0.14 M NH3 is titrated with 0.35 M HNO3.

a. Write a balance equation between NH3 and HNO3.

b. Predict the pKa of NH4+.

c. Determine the pH of the solution after the addition of 19.0 mL of HNO3.
Answer #1

as Balanced equation between NH₂ and HNO base HNO₃ + H₂ NHA + No, acid base a conjugate conjugate 5 mg of NH3 = 18 810*5 * (tin Hill hul fitration it is seen Hemaniw in eneessl amount that WHO tu Falcosa moles of why - (0,084 - 0.99.62) moles Xcess mso, por = -log [on] = -log (6.31x101) = 3.2 so, pH = 14- pOH = 14- 3-2 =